OpenStudy (samigupta8):

if x^2+px+1 is a factor of the expression ax^3+bx+c a.a^2+c^2=-ab b.a^2-c^2=-ab a^2-c^2=ab

OpenStudy (samigupta8):

@RadEn pls..hlp

OpenStudy (raden):

what's the question here ?

OpenStudy (samigupta8):

this is the question

OpenStudy (anonymous):

What is "the factor of the expression"? Sorry, my English is not so good.

OpenStudy (samigupta8):

ax^3+bx+c is any expression and x^2+px+1 is factor of it that's what ques says

OpenStudy (anonymous):

Thank you, I get it.

hartnn (hartnn):

lets try whats sum , product of roots for x^2+px+1 ? and whats for that cubic expression ?

OpenStudy (samigupta8):

sum is -p and product is 1

OpenStudy (samigupta8):

sum for cubic expression is 0 and product is c

hartnn (hartnn):

product is ? c? no...

OpenStudy (samigupta8):

bt why

hartnn (hartnn):

ok, so let 3 roots of cubic exp be A,B,C product of roots for cubic expression ax^3+bx^2+cx+d is -d/a , right ?

hartnn (hartnn):

so, here. won't it be -c/a ?

OpenStudy (samigupta8):

k...i undrstand it

hartnn (hartnn):

so, since x^2+px+1 is a factor of the expression ax^3+bx+c, two from A,B,C are also roots of x^2+px+1 got this ? this is important step

hartnn (hartnn):

let that 2 roots be A,B

OpenStudy (samigupta8):

kk...

hartnn (hartnn):

so, now we have AB =1 A+B=-p ABC=-c/a A+B+C =0 what about AB+BC+CA =...?

OpenStudy (samigupta8):

that's simply b

hartnn (hartnn):

not b/a ?

OpenStudy (samigupta8):

oh sorry i m considering a=1

OpenStudy (samigupta8):

then wt shud we do now

hartnn (hartnn):

don't. so, AB =1 A+B=-p ABC=-c/a A+B+C =0 AB+BC+CA =b/a everything clear till here ? see if you get C=p or not.....

OpenStudy (samigupta8):

yaa c=p

hartnn (hartnn):

also fro ABC = -c/a since AB=1 C=p=-c/a ok?

OpenStudy (samigupta8):

yaa i got it uptill here

hartnn (hartnn):

good. AB+BC+CA =b/a 1 + C (A+B) = b/a got this ? as AB=1

OpenStudy (samigupta8):

yaaaa

hartnn (hartnn):

now just plug in values! C = p = -b/a A+B = -p = b/a

OpenStudy (samigupta8):

c=p=-c/a then hw it came c=p=-b/a

hartnn (hartnn):

thats why i took capital c ---->C to avoid confusion C = p not c=p

OpenStudy (samigupta8):

no no i m saying dat C=p=-c/a and u r getting it as C=p=-b/a

hartnn (hartnn):

and my bad, typing mistake C = p = -c/a

hartnn (hartnn):

sorry

OpenStudy (samigupta8):

it's jst ok...

hartnn (hartnn):

C = p = -c/a A+B = -p = c/a

hartnn (hartnn):

1-c^2/a^2 = b/a ....